# Last Update: August, 30, 2024

### v1

@@
Last Update: August, 30, 2024
@@
Erzhi Liu111 erzhiliu@u.northwestern.edu. Center for Foundation Models and Generative AI & Department of Computer Science, Northwestern University, Evanston, IL 60208, USA  Jerry Yao-Chieh Hu222jhu@ensemblecore.ai; jhu@u.northwestern.edu. Ensemble AI, San Francisco, CA 94133, USA; Center for Foundation Models and Generative AI & Department of Computer Science, Northwestern University, Evanston, IL 60208, USA Alex Reneau333alex@ensemblecore.ai. Ensemble AI, San Francisco, CA 94133, USA  Zhao Song444 zsong@adobe.com; magic.linuxkde@gmail.com. Adobe Research, Seattle, WA 98103, USA  Han Liu555hanliu@northwestern.edu. Center for Foundation Models and Generative AI & Department of Computer Science & Department of Statistics and Data Science, Northwestern University, Evanston, IL 60208, USA
@@
We introduce a refined differentially private (DP) data structure for kernel density estimation (KDE), offering not only improved privacy-utility tradeoff but also better query efficiency over prior results. Specifically, we study the mathematical problem: given a similarity function $f$ (or DP KDE) and a private dataset $X\subset\mathbb{R}^{d}$, our goal is to preprocess $X$ so that for any query $y\in\mathbb{R}^{d}$, we approximate $\sum_{x\in X}f(x,y)$ in a differentially private fashion. The best previous algorithm for $f(x,y)=\|x-y\|_{1}$ is the node-contaminated balanced binary tree by [Backurs, Lin, Mahabadi, Silwal, and Tarnawski, ICLR 2024]. Their algorithm requires $O(nd)$ space and time for preprocessing with $n=|X|$. For any query point, the query time is $d\log n$, with an error guarantee of $(1+\alpha)$-approximation and $\epsilon^{-1}\alpha^{-0.5}d^{1.5}R\log^{1.5}n$.
@@ 1 Introduction
In this note, we focus on the $\ell_{1}$ kernel [^3] We remark that $\ell_{1}$ query distance is fundamental. Namely, it is easy to generalize the $\ell_{1}$ results to a wide range of kernels and distance functions following the provably dimensional reduction recipe of [BLM+24]. (i.e., $f(x,y)=\|x-y\|_{1}$). By far, the algorithm with the best privacy-utility tradeoff and query efficiency for Problem 1 is by Backurs, Lin, Mahabadi, Silwal, and Tarnawski [BLM+24]. They propose a DP data structure via a node-contaminated balanced binary tree for $\ell_{1}$ kernel. To be concrete, we begin by defining the similarity error between two data structures and then present their results.
@@
  • Query$(y\in\mathbb{R})$. It takes $y$ as input and spends $O(d\log n)$ time to output a scalar $A$ such that $\mathbb{E}[|A-A^{\prime}|]\leq O(\epsilon^{-1}Rd^{1.5}\log^{1.5}n)$.
@@ 2.4 High-Level Overview of Our DP Data Structure
    We make a key observation (Lemma 3.1) that these values ($s$ and $c$) are not depending on query $y$, while they are critical components of the final answer (the $\ell_{1}$ kernel $\sum_{x\in X}\|x-y\|_{1}$). This motivate us to construct the answer by using $L=\log n$ layers result, where each layer result is a smart combination of $y,s_{1},s_{2},c_{1},c_{2}$ (i.e, $(s_{1}-s_{2}+y\cdot c_{1}-y\cdot c_{2})$ ).
@@ 2.5 Related Works
However, despite their empirical [FC20, YLW23, WIL+23, NPF+24, ZWH+24, ZJL+23, JZLD21, TTE+23, SAT+23, MBA+23, HCL+24, LZL+24, EKB+24, MGA+24, PX23, GZCY23, ZNVA23, BLCZ22] and theoretical successes [BCW+23, WCWH23, KS24, ZL24, LWL+24, LWLC23, LIPO23, TLTO23, WHHL24, WHL+24, AS24, AS23, HSK+24, HYW+23, RSL+20, YBR+20, LLSS24, SWXL24, XSL24, LSSZ24b, LSSZ24a, LSSS24, LLS+24b, LLS+24a, LSSY24, WMS+24, CMFW24, FDG+24, FYWC24, WCL+24, DCWY24, GYY+24, HWSL24, LWCC24, ZLC23, WWY24, CCL+23a, CHZW23, OAS23, CCL+23b], there exists a gap from productizing many these advancements due to privacy, safety, and security concerns [QHZ+24, SHW+24, LHLW24, WMR+21]. These models, due to their vast scale and extensive training on large datasets, risk exposing sensitive information and amplifying biases present in the data. Privacy issues arise when models inadvertently memorize and reproduce personal data from training sets [CHN+23, CTW+21]. Safety concerns include the potential for generating harmful or misleading content, especially in applications where accuracy is paramount [WMR+21, Ope23, TAB+23, Ant23, YLH+24, LYZ+24, SCB+24, DLL+23, LDX+23, LXCX23, SPT+23, YLX23]. Security vulnerabilities also exist, as these models may be susceptible to adversarial attacks that manipulate outputs or extract proprietary information [ZYZ+24, JXN+24, XQZ+24]. Addressing these challenges is essential to ensure the responsible and secure use of foundation models across various applications. In this work, we focus on the privacy of foundation models at a fundamental level by study the formalized KDE query Problem 1.
@@ 3.2 Time Complexity
If he total layer $L=\log(n)$, the running time of Init (Algorithm 1) is $O(n)$.
@@ 3.4 Error Guarantee
3.5 One Dimensional Data Structure
Theorem 3.1.
Given a dataset $X\subset\mathbb{R}$ with $|X|$. There is an algorithm that uses $O(n)$ space to build a data-structure (Algorithm 1, 2) which supports the following operations
  • Init$(X,\epsilon)$ It takes dataset $X$ and privacy parameter $\epsilon$ as input and spends $O(n)$ time to build a data-structure.
@@ 3.6 High Dimensional $\ell_{1}$ Distance Query
By applying our one dimensional data structure on each dimension independently, we calculate high dimensional queries.
@@
Let $X=\{x_{i}\}_{i\in[n]}$ be a dataset of $n$ $d$ dimensional vectors $x_{i}\in[0,R)^{d}$. Let $A^{\prime}=\sum_{i=1}^{n}|x_{i}-y|_{1}$ be the true distance query value. Let $A$ be the output of Query (Algorithm 3). Then we have $\mathbb{E}[|A-A^{\prime}|]\leq O(\epsilon^{-1}Rd^{1.5}\log^{1.5}n)$.
@@
Theorem 3.2 (Formal version of Theorem 1.1).
Given a dataset $X\subset\mathbb{R}^{d}$ with $|X|$. There is an algorithm that uses $O(nd)$ space to build a data-structure (Algorithm 1, 2) which supports the following operations
  • Init$(X\subset\mathbb{R}^{d},\epsilon\in(0,1))$. It takes dataset $X$ and privacy parameter $\epsilon\in(0,1)$ as input and spends $O(nd)$ time to build a data-structure.
@@
By Lemma 3.6, we prove the privacy part of the Theorem.

### v2

# Last Update: September, 5, 2024

@@
Last Update: September, 5, 2024
@@
Erzhi Liu111 erzhiliu@u.northwestern.edu. Center for Foundation Models and Generative AI & Department of Computer Science, Northwestern University, Evanston, IL 60208, USA  Jerry Yao-Chieh Hu222jhu@ensemblecore.ai; jhu@u.northwestern.edu. Ensemble AI, San Francisco, CA 94133, USA; Center for Foundation Models and Generative AI & Department of Computer Science, Northwestern University, Evanston, IL 60208, USA Alex Reneau333alex@ensemblecore.ai. Ensemble AI, San Francisco, CA 94133, USA  Zhao Song444magic.linuxkde@gmail.com. Simons Institute for the Theory of Computing, UC Berkeley, Berkeley, CA 94720, USA  Han Liu555hanliu@northwestern.edu. Center for Foundation Models and Generative AI & Department of Computer Science & Department of Statistics and Data Science, Northwestern University, Evanston, IL 60208, USA
@@
We introduce a refined differentially private (DP) data structure for kernel density estimation (KDE) with $\ell_{1}$, $\ell_{2}$ and $\ell_{p}^{p}$ kernels. This new DP data structure offers not only improved privacy-utility tradeoff but also better query efficiency over prior results. Specifically, we study the mathematical problem: given a similarity function $f$ (or DP KDE) and a private dataset $X\subset\mathbb{R}^{d}$, our goal is to preprocess $X$ so that for any query $y\in\mathbb{R}^{d}$, we approximate $\sum_{x\in X}f(x,y)$ in a differentially private fashion. The best previous algorithm for $f(x,y)=\|x-y\|_{1}$ is the node-contaminated balanced binary tree by [Backurs, Lin, Mahabadi, Silwal, and Tarnawski, ICLR 2024]. Their algorithm requires $O(nd)$ space and time for preprocessing with $n=|X|$. For any query point, the query time is $d\log n$, with an error guarantee of $(1+\alpha)$-approximation and $\epsilon^{-1}\alpha^{-0.5}d^{1.5}R\log^{1.5}n$.
@@ 1 Introduction
In this note, we focus on the $\ell_{1}$ kernel [^3] We remark that $\ell_{1}$ query distance is fundamental. Namely, it is easy to generalize the $\ell_{1}$ results to a wide range of kernels and distance functions following the provably dimensional reduction recipe of [BLM+24]. See Appendix B and C for improved DP KDE results with $\ell_{2}$ and $\ell_{p}^{p}$ kernels. (i.e., $f(x,y)=\|x-y\|_{1}$). By far, the algorithm with the best privacy-utility tradeoff and query efficiency for Problem 1 is by Backurs, Lin, Mahabadi, Silwal, and Tarnawski [BLM+24]. They propose a DP data structure via a node-contaminated balanced binary tree for $\ell_{1}$ kernel. To be concrete, we begin by defining the similarity error between two data structures and then present their results.
@@
  • Query$(y\in\mathbb{R}^{d})$. It takes $y$ as input and spends $O(d\log n)$ time to output a scalar $A$ such that $\mathbb{E}[|A-A^{\prime}|]\leq O(\epsilon^{-1}Rd^{1.5}\log^{1.5}n)$.
@@
Lastly, we remark that our $\ell_{1}$ results are transferable to other kernels via the provably dimensional reduction recipe of [BLM+24]. To showcase this, we generalize our results to $\ell_{2}$ and $\ell_{p}^{p}$ kernels following [BLM+24]. Notably, these results also improve upon the best previous results reported in [BLM+24] for $\ell_{2}$ and $\ell_{p}^{p}$ with $p\in[1,n]$. For details, see Appendix B and C, particularly the comparison Tables 2 and 3.
@@ 2.4 High-Level Overview of Our DP Data Structure
    We make a key observation (Lemma 3.1) that these values ($s$ and $c$) are not depending on query $y$, while they are critical components of the final answer (the $\ell_{1}$ kernel $\sum_{x\in X}\|x-y\|_{1}$). This motivate us to construct the answer by using $L=\log n$ layers result, where each layer result is a smart combination of $y,s_{1},s_{2},c_{1},c_{2}$ (i.e., $(s_{1}-s_{2}+y\cdot c_{1}-y\cdot c_{2})$ ).
@@ 2.5 Related Works
However, despite their empirical [FC20, YLW23, WIL+23, NPF+24, ZWH+24, ZJL+23, JZLD21, TTE+23, SAT+23, MBA+23, HCL+24, LZL+24, EKB+24, MGA+24, PX23, GZCY23, ZNVA23, BLCZ22] and theoretical successes [BCW+23, WCWH23, KS24, ZL24, LWL+24, LWLC23, LIPO23, TLTO23, WHHL24, WHL+24, AS24, AS23, HSK+24, HYW+23, RSL+20, YBR+20, LLSS24, SWXL24, XSL24, LSSZ24b, LSSZ24a, LSSS24, LLS+24b, LLS+24a, LSSY24, WMS+24, CMFW24, FDG+24, FYWC24, WCL+24, DCWY24, GYY+24, HWL+24, LWCC24, ZLC23, WWY24, CCL+23a, CHZW23, OAS23, CCL+23b], there exists a gap from productizing many these advancements due to privacy, safety, and security concerns [QHZ+24, SHW+24, LHLW24, WMR+21]. These models, due to their vast scale and extensive training on large datasets, risk exposing sensitive information and amplifying biases present in the data. Privacy issues arise when models inadvertently memorize and reproduce personal data from training sets [CHN+23, CTW+21]. Safety concerns include the potential for generating harmful or misleading content, especially in applications where accuracy is paramount [WMR+21, Ope23, TAB+23, Ant23, YLH+24, LYZ+24, SCB+24, DLL+23, LDX+23, LXCX23, SPT+23, YLX23]. Security vulnerabilities also exist, as these models may be susceptible to adversarial attacks that manipulate outputs or extract proprietary information [ZYZ+24, JXN+24, XQZ+24]. Addressing these challenges is essential to ensure the responsible and secure use of foundation models across various applications. In this work, we focus on the privacy of foundation models at a fundamental level by study the formalized KDE query Problem 1.
@@ 3.2 Time Complexity
Here we compute the init and query time of Algorithm 1 and 2, respectively.
@@
If the total layer $L=\log(n)$, the running time of Init (Algorithm 1) is $O(n)$.
@@ 3.3 Privacy Guarantees
Here we provide our DP guarantee of our proposed DP data structure (i.e., Algorithm 1).
@@ 3.4 Error Guarantee
Here we provide the (data-structure) similarity error of our proposed DP data structure.
@@
3.5 One Dimensional Differentially Private Data Structure
Collectively, we present the init time, space, query time, approximation error, and privacy guarantees for the 1-dimensional PD data structure (Algorithm 1, 2) in the next theorem. As a reminder, we address the 1-dimensional KDE query problem (Problem 1 with $d=1$) using the $\ell_{1}$ kernel distance.
Theorem 3.1 ($1$-Dimensional $\ell_{1}$ DP KDE Query).
Given a dataset $X\subset\mathbb{R}$ with $|X|=n$, there is an algorithm that uses $O(n)$ space to build a data structure (Algorithm 1, 2) which supports the following operations
  • Init$(X,\epsilon)$. It takes dataset $X$ and privacy parameter $\epsilon$ as input and spends $O(n)$ time to build a data-structure.
@@ 3.6 High Dimensional $\ell_{1}$ Distance Query
We now generalize our 1-dimensional DP data structure to $d$-dimensional DP KDE queries. Specifically, we apply our 1-dimensional DP data structure independently to each dimension to compute high-dimensional queries.
@@
Let $X=\{x_{i}\}_{i\in[n]}$ be a dataset of $n$ $d$-dimensional vectors $x_{i}\in[0,R)^{d}$. Let $A^{\prime}=\sum_{i=1}^{n}|x_{i}-y|_{1}$ be the true distance query value. Let $A$ be the output of Query (Algorithm 3). Then we have $\mathbb{E}[|A-A^{\prime}|]\leq O(\epsilon^{-1}Rd^{1.5}\log^{1.5}n)$.
@@
Theorem 3.2 ($d$-Dimensional $\ell_{1}$ DP KDE Query, Formal version of Theorem 1.1).
Given a dataset $X\subset\mathbb{R}^{d}$ with $|X|=n$, there is an algorithm that uses $O(nd)$ space to build a data-structure (Algorithm 1, 2) which supports the following operations
  • Init$(X\subset\mathbb{R}^{d},\epsilon\in(0,1))$. It takes dataset $X$ and privacy parameter $\epsilon\in(0,1)$ as input and spends $O(nd)$ time to build a data structure.
@@
By Lemma 3.6, we prove the privacy guarantees.